Consider an asset that costs $519,200 and is depreciated straight-line to zero over its 7-year tax life. The asset is to be used in a 3-year project; at the end of the project, the asset can be sold for $64,900.

If the relevant tax rate is 34 percent, what is the aftertax cash flow from the sale of this asset? (Do not round your intermediate calculations.)

A. $150,892.50

B. $143,707.14

C. $136,521.78

D. $42,834.00

E. $793,090.00
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(519200/7)=$74171.42857/year

Hence book value as on date of sale=cost-accumulated depreciation

=519200-(74171.42857*3)

=$296685.7143

Hence loss on sale =$296685.7143-$64900

=$231785.7143

Hence aftertax cash flow=Sale proceeds+(Tax rate*loss on sale)

=64900+($231785.7143*0.34)

which is equal to

=$143,707.14(Approx).